OpenStudy (anonymous):

Is there a method for finding the derivative of y=x^2 without using calculus OR LIMITS? I think I have found a way, but would first like to hear if you have heard of any.

OpenStudy (anonymous):

did you take two half derivatives?

OpenStudy (anonymous):

I think not (don't know what it is)

OpenStudy (anonymous):

If you found another way, you probably did calculus without realizing it..

OpenStudy (anonymous):

No, definitely not

Parth (parthkohli):

Power Rule? Well, that is Calculus too.

OpenStudy (anonymous):

No differentiation rules at all

OpenStudy (anonymous):

The same method can be applied to y=1/x

OpenStudy (anonymous):

Ok, well if no-one else has anything to say, I'll proceed

OpenStudy (anonymous):

you have an audience now. may we see the method?

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

Let D(0,y1) be a point on the y-axis We want to find the equation of a line such that it passes through the point D and is a tangent to the parabola y=x^2. The equation can be written in the form: y=mx+y1 To find where the line and parabola contact, solve simultaneously. x^2=mx+y1 x^2-mx-y1=0 x= -b±√∆ ------ 2a but since it is a tangent, only one point of contact. So ∆=0 x = m/2 m=2x Tada

OpenStudy (anonymous):

Does your method generalize? (past quadratic)

OpenStudy (anonymous):

It might, so far only works for x^2 and 1/x

OpenStudy (anonymous):

x^2=mx+y1 x^2-mx-y1=0 Aren't these the same?

OpenStudy (anonymous):

yes, just rearranging to show clearly what the a,b,c values are

OpenStudy (anonymous):

@estudier Try it for 1/x

OpenStudy (anonymous):

I will, just trying to get it first...

OpenStudy (anonymous):

This is no calculus, right?

OpenStudy (anonymous):